. Find the nature of the roots of the following quadratic equations. If the real roots exist, find them; (i) 2x 2 - 3x + 5 = 0

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,
2x^2 - 3x + 5 = 0
Comparing the equation with ax^2 + bx + c = 0,
we get a = 2, b = -3 and c = 5
We know,
Discriminant = b^2 - 4ac
= ( - 3)^2 - 4 (2) (5)
= 9 - 40 = - 31
As you can see, b^2 - 4ac < 0
Therefore, no real root is possible for the given equation.

Thanks